10ml of Hydrochloric acid solution of pH=1 is mixed with 40ml of a hydrochloric acid solution of pH=4. The pH of the mixture is?


1
Expert's answer
2021-07-21T04:37:02-0400


According to the given question HCL+NaOH+NaCL+H2O

As per details that are given in the question we can calculate the following

No. Of moles of HCL= (40ml) (.1mol/L)

= (0.04L) (0.1mol/L) = 0.004 moles

No. Of moles of NaOH = (10ml) (.45moles/L)

= (0.01L) (0.45mol/L) = 0.0045 moles

Therefore, excess amount of NaOH will be = 0.0045moles - 0.004moles = 0.0005 moles

Therefore the total volume will be

= 0.04 + 0.01 = 0.05L

Molarity of OH will be = 0.0005moles/0.05L= 0.01M

pOH = -log (OH)

= - log 0.01

Therefor pOH = 2

So pH = 14 - pOH

14 - 2 = 12